use exact values thank you
sin (arcsin to arccos e/w?
Answer #1


are cos(83) Sin ( are sint are sin(t) = sint (1/2) sint (sna) z Sin ( 24 = sin ( Ga-un) ។ au areas (83) = 65 (V3,,) - (est) -